# A Bombieri-Vinogradov theorem for sectors in real quadratic number fields

### v1

@@ 1.1. History
One way to get around this is to replace congruences $a\equiv b\bmod{m}$ of integers by ray class equivalences $\mathfrak{a}\sim\mathfrak{b}\bmod{\mathfrak{m}}$ of ideals modulo a modulus $\mathfrak{m}$ for $\mathbb{K}$, as introduced by Hecke. A generalized Bombieri-Vinogradov type theorem of this kind was established by Wilson [19] and improved by Huxley [9]. One may also proceed in a different direction if $\mathbb{K}$ is Galois over $\mathbb{Q}$: Instead of prime ideals in ray classes, one may count rational primes $p\equiv a\bmod{q}$ which are unramified in $\mathbb{K}$ and whose Artin symbol $(p,\mathbb{K}/\mathbb{Q})$ equals a given conjugacy class $C$ in the Galois group of $\mathbb{K}$ over $\mathbb{Q}$. This route was taken by Murty and Murty [15] and later generalized to the situation of non-Galois extensions $\mathbb{K}$ of a general base field by Murty and Peterson [16]. The congruence condition above can be expressed in the form $\mathcal{N}(\mathfrak{p})\equiv a\bmod{q}$, where $\mathcal{N}(\mathfrak{p})$ is the norm of a prime ideal $\mathfrak{p}$ in $\mathcal{O}_{\mathbb{K}}$. The above results in [15] and [16] were strengthened by Jiang, Lü and Wang [10] and refined by Thorner [18] who included a short interval condition. Khale, O’Kuhn, Panidapu, Sun and Zhang [12] considered versions for Galois extensions of $\mathbb{Q}$ with a small sector condition determined by Hecke characters. However, also in this setting, the congruence condition was of the form $\mathcal{N}(\mathfrak{p})\equiv a\bmod{q}$.
Hinz [8] developed a version of the Bombieri-Vinogradov theorem in which he considered a genuine congruence condition of the form $p\equiv a\bmod{\mathfrak{q}}$ in $\mathcal{O}_{\mathbb{K}}$. So he really counted prime elements $p\in\mathcal{O}_{\mathbb{K}}$ in residue classes $a\bmod{\mathfrak{q}}$ such that the norm of $p$ over $\mathbb{Q}$ is bounded by $x$. For this to work, he restricted these prime elements $p$ to a box. More in detail, his restriction was that $\left(p_{1},p_{2},...,p_{r_{1}},|p_{r_{1}+1}|^{2},|p_{r_{1}+2}|^{2},...,|p_{r_{1}+r_{2}}|^{2}\right)$ lies in a box $(0,y_{1}]\times(0,y_{r_{1}+r_{2}}]$, where $p_{1},...,p_{r_{1}}$ are the real and $p_{r_{1}+1},\overline{p_{r_{1}+1}},...,p_{r_{1}+r_{2}},\overline{p_{r_{1}+r_{2}}}$ the complex embeddings of $p$. He assumed that the $y_{k}$’s are about of size $x^{1/(r_{1}+r_{2})}$ so that the volume of the box is about of size $x$.
@@
1.2. General notations
To state our main result, we first set up some notations. For a number field $\mathbb{K}$, we denote by $\mathcal{O}_{\mathbb{K}}$ its ring of algebraic integers, by $\mathcal{O}_{\mathbb{K}}^{\ast}$ the group of units in $\mathcal{O}_{\mathbb{K}}$ and by $\mathcal{I}_{\mathbb{K}}$ the set of non-zero integral ideals in $\mathcal{O}_{\mathbb{K}}$. For $\mathfrak{a},\mathfrak{b}\in\mathcal{I}_{\mathbb{K}}$, we write $(\mathfrak{a},\mathfrak{b})=1$ if $\mathfrak{a}$ and $\mathfrak{b}$ have no common prime ideal divisors. For $a\in\mathcal{O}_{\mathbb{K}}$ and $\mathfrak{a}\in\mathcal{I}_{\mathbb{K}}$, we write $(a,\mathfrak{a})=1$ iff $((a),\mathfrak{a})=1$. If $\mathfrak{a}\in\mathcal{I}_{\mathbb{K}}$, then we define the norm of the ideal $\mathfrak{a}$ by $\mathcal{N}(\mathfrak{a}):=[\mathcal{O}_{\mathbb{K}}:\mathfrak{a}]$. If $a\in\mathbb{K}$, then we set $\mathcal{N}(a):=|N_{\mathbb{K}:\mathbb{Q}}(a)|$. We note that $\mathcal{N}(a)=\mathcal{N}((a))$ if $a\in\mathcal{O}_{\mathbb{K}}$. We define the Von Mangoldt function on $\mathcal{I}_{\mathbb{K}}$ by
@@ 1.2. General notations
From now onward, we restrict $\mathbb{K}$ to be a real quadratic number field, i.e. $\mathbb{K}=\mathbb{Q}(\sqrt{d})$, where $d>1$ is a square-free integer. Throughout the sequel, for simplicity, we shall also assume that $\mathbb{K}$ has class number one, i.e. $\mathcal{O}_{\mathbb{K}}$ is a principal ideal domain. We denote the two embeddings of $\mathbb{K}$, the identity and conjugation, by
@@
We will use the following lemma to suitably restrict generators of (principal) ideals.
@@
Let $\mathfrak{a}\in{I}_{\mathbb{K}}$. Then there exists a generator $a$ of $\mathfrak{a}$, unique up to the sign, such that
@@
where $\eta_{1}$ and $\eta_{2}$ are fixed such that $1\leq\eta_{1}<\eta_{2}\leq 1$. By Lemma 1, every $\mathfrak{a}\in\mathcal{I}_{\mathbb{K}}$ has exactly one generator in S(1,1)S(-1,1). Throughout the following, let $x\geq 2$. For $a\in\mathcal{O}_{\mathbb{K}}$ and $\mathfrak{q}\in\mathcal{I}_{\mathbb{K}}$ with $(a,\mathfrak{q})=1$, we set
@@ 1.3. Main result
Let $\mathbb{K}$ be a real quadratic field of class number one. Let $x\geq 2$, $1\leq Q\leq x^{2/3}$ and $A>0$. Keep the notations above. Then
(1.5) $\sum\limits_{\begin{subarray}{c}\mathfrak{q}\in\mathcal{I}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{q})\leq Q\end{subarray}}\,\,\max\limits_{\begin{subarray}{c}a\bmod\mathfrak{q}\\ (a,\mathfrak{q})=1\end{subarray}}\,\,\max\limits_{y\leq x}\left|E_{S}(y;\mathfrak{q},a)\right|\ll_{A,\eta,\mathbb{K}}\left(x^{13/15}Q^{8/15}+x^{2/3}Q^{5/3}\right)\log^{11}x+\frac{x}{\log^{A}x}.$
In particular, if $Q\leq x^{1/5}(\log x)^{-3(A+11)/5}$, we have
$\sum\limits_{\begin{subarray}{c}\mathfrak{q}\in\mathcal{I}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{q})\leq Q\end{subarray}}\,\,\max\limits_{\begin{subarray}{c}a\bmod\mathfrak{q}\\ (a,\mathfrak{q})=1\end{subarray}}\,\,\max\limits_{y\leq x}\left|E_{S}(y;\mathfrak{q},a)\right|\ll_{A,\eta,\mathbb{K}}\frac{x}{\log^{A}x}.$
Remark 1. Using Lemma 1 and Landau’s prime ideal theorem for number fields, we may further approximate the main term $M_{S}(x;\mathfrak{q},a)$ by
$\begin{split}M_{S}(x;\mathfrak{q},a)=\frac{\eta}{\varphi(\mathfrak{q})}\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{I}_{\mathbb{K}}\\ \ \mathcal{N}(\mathfrak{s})\leq x\\ (\mathfrak{s},\mathfrak{q})=1\end{subarray}}\Lambda(\mathfrak{s})=&\frac{\eta}{\varphi(\mathfrak{q})}\Bigg{(}\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{I}_{\mathbb{K}}\\ \ \mathcal{N}(\mathfrak{s})\leq x\end{subarray}}\Lambda(\mathfrak{s})+O\left(\log^{2}(\mathcal{N}(\mathfrak{q}x))\right)\Bigg{)}\\ =&\frac{\eta}{\varphi(\mathfrak{q})}\left(x+O_{A}\left(\frac{x}{(\log x)^{A}}\right)\right),\end{split}$
provided that $\mathcal{N}(\mathfrak{q})\leq x$. Therefore, the estimates in Theorem 1 remain valid if the term $E_{S}(y;\mathfrak{q},a)$ is replaced by
$E_{S}^{\prime}(y;\mathfrak{q},a):=\psi_{S}(y;\mathfrak{q},a)-\frac{\eta}{\varphi(\mathfrak{q})}\cdot x.$
@@ 1.4. Organization of this article
After introducing the necessary tools, we will use Fourier analysis to pick out the residue class and sector conditions via Hecke characters. This reduces the problem to bounding averages of Hecke character sums over prime ideals (in fact, we work with the von Mangoldt function instead of the indicator function of primes). The moduli $\mathfrak{q}$ with a small norm will be treated using a version of the Siegel-Walfisz theorem. For the remaining contribution, we will use Vaughan’s identity to reduce the problem to bounding averages of bilinear sums with Hecke characters over ideals. For the estimation of these averages, we will use a version of the Polya-Vinogradov inequality as well as a large sieve inequality for Hecke characters. This will finally allow us to establish our Bombieri-Vinogradov theorem for sectors in real quadratic fields. Throughout the remainder of this article, we allow implied constants to depend on $\mathbb{K}$ and $\eta$.
Remark 2: The somewhat disappointing level of distribution of $1/5-\varepsilon$ in our Theorem 1 arises from a large sieve inequality for a family of Hecke characters (Lemma 6 below) which is worse than what we expect to hold (see Remark 3 after the proof of Lemma 6). It would be desirable to improve this bound. Our conjectural large sieve inequality would lead to a level of distribution $1/2-\varepsilon$. Another future task could be to remove the class number one condition, which shouldn’t be hard.
@@ 2.1. Hecke characters for real quadratic fields
In this subsection, we provide an explicit description of Hecke characters for real quadratic fields, based on our more detailed description in [2, section 2], which in turn was based on material in [6], [7], [14, section 3.3], and [17]. Hecke characters are characters of the group of fractional ideals of $\mathbb{K}$ which are coprime to some modulus. We will use them to detect elements of $\mathcal{O}_{\mathbb{K}}$ in intersections of sectors and arithmetic progressions. To describe them properly, we set up some notations below.
@@
when $\alpha\in\mathbb{K}$ is such that $(\alpha)$ is coprime to $\mathfrak{q}$. If $(\alpha)$ is not coprime to $\mathfrak{q}$, then we set $\chi(\alpha):=0$. It is easy to see that $\chi(\alpha)$ is well-defined. We refer to $\chi$ as a Dirichlet character modulo $\mathfrak{q}$ as well. On the principal fractional ideals of $\mathbb{K}$ (which are the only fractional ideals since we assumed $\mathbb{K}$ to have class number one), Hecke characters modulo $\mathfrak{q}$ take the form
@@
where $\chi$ is a Dirichlet character modulo $\mathfrak{q}$ and $\chi_{\infty}$ is referred to as infinite part of $\chi_{\text{Hecke}}$. We say that the Hecke character $\chi_{\text{Hecke}}$ belongs to the Dirichlet character $\chi$. (More precisely, $\chi_{\text{Hecke}}$ should be viewed as a Hecke character to a modulus with finite part $\mathfrak{q}$ and infinite part $\mathfrak{m}_{\infty}$. For simplicity, we surpress $\mathfrak{m}_{\infty}$ throughout this article.) For $\chi_{\text{Hecke}}((\alpha))$ to be well-defined, it is necessary that
@@ 2.2. Sums with Hecke Characters
Let $X\geq 1$ and $\mathbb{K}$ be a number field of degree $n$ over $\mathbb{Q}$. Let $\chi_{\text{\rm Hecke}}$ be a non-trivial Hecke character modulo an ideal $\mathfrak{q}\in\mathcal{I}_{\mathbb{K}}$. Then
@@
Let $X\geq 2$ and $\mathbb{K}$ be a number field and $A>0$. Then there exists a positive constant $C=C(A)$ such that for any primitive Hecke character $\chi_{\text{\rm Hecke}}$ modulo an ideal $\mathfrak{q}\in\mathcal{I}_{\mathbb{K}}$ with $\mathcal{N}(\mathfrak{q})\leq\log^{A}X$, we have

### v2

@@ 1.1. History
One way to get around this is to replace congruences $a\equiv b\bmod{m}$ of integers by ray class equivalences $\mathfrak{a}\sim\mathfrak{b}\bmod{\mathfrak{m}}$ of ideals to a modulus $\mathfrak{m}$ for $\mathbb{K}$, as introduced by Hecke. A generalized Bombieri-Vinogradov type theorem of this kind was established by Wilson [19] and improved by Huxley [9]. One may also proceed in a different direction if $\mathbb{K}$ is Galois over $\mathbb{Q}$: Instead of prime ideals in ray classes, one may count rational primes $p\equiv a\bmod{q}$ which are unramified in $\mathbb{K}$ and whose Artin symbol $(p,\mathbb{K}/\mathbb{Q})$ equals a given conjugacy class $C$ in the Galois group of $\mathbb{K}$ over $\mathbb{Q}$. This route was taken by Murty and Murty [15] and later generalized to the situation of non-Galois extensions $\mathbb{K}$ of a general base field by Murty and Peterson [16]. The congruence condition above can be expressed in the form $\mathcal{N}(\mathfrak{p})\equiv a\bmod{q}$, where $\mathcal{N}(\mathfrak{p})$ is the norm of a prime ideal $\mathfrak{p}$ in $\mathcal{O}_{\mathbb{K}}$. The above results in [15] and [16] were strengthened by Jiang, Lü and Wang [10] and refined by Thorner [18] who included a short interval condition. Khale, O’Kuhn, Panidapu, Sun and Zhang [12] considered versions for Galois extensions of $\mathbb{Q}$ with a small sector condition determined by Hecke characters. However, also in this setting, the congruence condition was of the form $\mathcal{N}(\mathfrak{p})\equiv a\bmod{q}$.
Hinz [8] developed a version of the Bombieri-Vinogradov theorem in which he considered a genuine congruence condition of the form $p\equiv a\bmod{\mathfrak{q}}$ in $\mathcal{O}_{\mathbb{K}}$. So he really counted prime elements $p\in\mathcal{O}_{\mathbb{K}}$ in residue classes $a\bmod{\mathfrak{q}}$ such that the norm of $p$ over $\mathbb{Q}$ is bounded by $x$. For this to work, he restricted these prime elements $p$ to a box. More in detail, his restriction was that $\left(p_{1},p_{2},...,p_{r_{1}},|p_{r_{1}+1}|^{2},|p_{r_{1}+2}|^{2},...,|p_{r_{1}+r_{2}}|^{2}\right)$ lies in a box $(0,y_{1}]\times\cdots\times(0,y_{r_{1}+r_{2}}]$, where $p_{1},...,p_{r_{1}}$ are the real and $p_{r_{1}+1},\overline{p_{r_{1}+1}},...,p_{r_{1}+r_{2}},\overline{p_{r_{1}+r_{2}}}$ the complex embeddings of $p$. He assumed that the $y_{k}$’s are about of size $x^{1/(r_{1}+r_{2})}$ so that the volume of the box is about of size $x$.
@@
1.2. General notation
To state our main result, we first set up some notation. For a number field $\mathbb{K}$, we denote by $h$ its class number, by $\mathcal{O}_{\mathbb{K}}$ its ring of algebraic integers, by $\mathcal{O}_{\mathbb{K}}^{\ast}$ the group of units in $\mathcal{O}_{\mathbb{K}}$, by $\mathcal{I}_{\mathbb{K}}$ the set of non-zero integral ideals in $\mathcal{O}_{\mathbb{K}}$, by $\mathcal{P}_{\mathbb{K}}$ the subset of principal ideals, by $\mathcal{I}^{\text{frac}}_{\mathbb{K}}$ the group of fractional ideals and by $\mathcal{P}^{\text{frac}}_{\mathbb{K}}$ the subgroup of principal fractional ideals. For $\mathfrak{a},\mathfrak{b}\in\mathcal{I}_{\mathbb{K}}$, we write $(\mathfrak{a},\mathfrak{b})=1$ if $\mathfrak{a}$ and $\mathfrak{b}$ have no common prime ideal divisors. For $a\in\mathcal{O}_{\mathbb{K}}$ and $\mathfrak{a}\in\mathcal{I}_{\mathbb{K}}$, we write $(a,\mathfrak{a})=1$ iff $((a),\mathfrak{a})=1$. If $\mathfrak{a}\in\mathcal{I}_{\mathbb{K}}$, then we define the norm of the ideal $\mathfrak{a}$ by $\mathcal{N}(\mathfrak{a}):=[\mathcal{O}_{\mathbb{K}}:\mathfrak{a}]$. If $a\in\mathbb{K}$, then we set $\mathcal{N}(a):=|N_{\mathbb{K}:\mathbb{Q}}(a)|$. We note that $\mathcal{N}(a)=\mathcal{N}((a))$ if $a\in\mathcal{O}_{\mathbb{K}}$. We define the von Mangoldt function on $\mathcal{I}_{\mathbb{K}}$ by
@@ 1.2. General notation
From now onward, we restrict $\mathbb{K}$ to be a real quadratic number field, i.e. $\mathbb{K}=\mathbb{Q}(\sqrt{d})$, where $d>1$ is a square-free integer. We denote the two embeddings of $\mathbb{K}$, the identity and conjugation, by
@@
We will use the following lemma to suitably restrict generators of principal ideals.
@@
Let $\mathfrak{a}\in\mathcal{P}_{\mathbb{K}}$. Then there exists a generator $a$ of $\mathfrak{a}$, unique up to the sign, such that
@@
where $\eta_{1}$ and $\eta_{2}$ are fixed such that $1\leq\eta_{1}<\eta_{2}\leq 1$. By Lemma 1, every $\mathfrak{a}\in\mathcal{P}_{\mathbb{K}}$ has exactly one generator in S(1,1)S(-1,1). Throughout the following, let $x\geq 2$. For $a\in\mathcal{O}_{\mathbb{K}}$ and $\mathfrak{q}\in\mathcal{I}_{\mathbb{K}}$ with $(a,\mathfrak{q})=1$, we set
@@ 1.3. Main result
Let $\mathbb{K}$ be a real quadratic field. Let $x\geq 2$, $1\leq Q\leq x^{4/5}$, $\varepsilon>0$ and $A>0$. Keep the notations above. Then
(1.5) $\sum\limits_{\begin{subarray}{c}\mathfrak{q}\in\mathcal{I}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{q})\leq Q\end{subarray}}\,\,\max\limits_{\begin{subarray}{c}a\bmod\mathfrak{q}\\ (a,\mathfrak{q})=1\end{subarray}}\,\,\max\limits_{y\leq x}\left|E_{S}(y;\mathfrak{q},a)\right|\ll_{\varepsilon,A,\eta,\mathbb{K}}\left(x^{9/11}Q^{8/11}+x^{1/2}Q^{2}\right)x^{\varepsilon}+\frac{x}{\log^{A}x}.$
In particular, if $Q\leq x^{1/4-2\varepsilon}$, we have
$\sum\limits_{\begin{subarray}{c}\mathfrak{q}\in\mathcal{I}_{\mathbb{K}}\\ \mathcal{N}(\mathfrak{q})\leq Q\end{subarray}}\,\,\max\limits_{\begin{subarray}{c}a\bmod\mathfrak{q}\\ (a,\mathfrak{q})=1\end{subarray}}\,\,\max\limits_{y\leq x}\left|E_{S}(y;\mathfrak{q},a)\right|\ll_{\varepsilon,A,\eta,\mathbb{K}}\frac{x}{\log^{A}x}.$
Remark 1. Using Lemma 1 and the Chebotarev density theorem, we may further approximate the main term $M_{S}(x;\mathfrak{q},a)$ by
$\begin{split}M_{S}(x;\mathfrak{q},a)=\frac{\eta}{\varphi(\mathfrak{q})}\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{P}_{\mathbb{K}}\\ \ \mathcal{N}(\mathfrak{s})\leq x\\ (\mathfrak{s},\mathfrak{q})=1\end{subarray}}\Lambda(\mathfrak{s})=&\frac{\eta}{\varphi(\mathfrak{q})}\Bigg(\sum\limits_{\begin{subarray}{c}\mathfrak{s}\in\mathcal{P}_{\mathbb{K}}\\ \ \mathcal{N}(\mathfrak{s})\leq x\end{subarray}}\Lambda(\mathfrak{s})+O\left(\log^{2}(\mathcal{N}(\mathfrak{q}x))\right)\Bigg)\\ =&\frac{\eta}{h\varphi(\mathfrak{q})}\left(x+O\left(\frac{x}{(\log x)^{A}}\right)\right)\end{split}$
for any $A>0$, provided that $\mathcal{N}(\mathfrak{q})\leq x$. Therefore, the estimates in Theorem 1 remain valid if the term $E_{S}(y;\mathfrak{q},a)$ is replaced by
$E_{S}^{\prime}(y;\mathfrak{q},a):=\psi_{S}(y;\mathfrak{q},a)-\frac{\eta}{h\varphi(\mathfrak{q})}\cdot x.$
@@ 1.4. Organization of this article
After introducing the necessary tools, we will use Fourier analysis to pick out the residue class and sector conditions via Hecke characters. This reduces the problem to bounding averages of Hecke character sums over prime ideals (in fact, we work with the von Mangoldt function instead of the indicator function of primes). The moduli $\mathfrak{q}$ with a small norm will be treated using a version of the Siegel-Walfisz theorem. For the remaining contribution, we will use Vaughan’s identity to reduce the problem to bounding averages of bilinear sums with Hecke characters over ideals. For the estimation of these averages, we will use a version of the Polya-Vinogradov inequality as well as a large sieve inequality for Hecke characters. This will finally allow us to establish our Bombieri-Vinogradov theorem for sectors in real quadratic fields. Throughout the remainder of this article, we allow implied constants to depend on $\varepsilon$, $A$, $\eta$ and $\mathbb{K}$.
Remark 2: The level of distribution of $1/4-\varepsilon$ in our Theorem 1 arises from a large sieve inequality for a family of Hecke characters (Lemma 6 below) which is worse than what we expect to hold (see Remark 3 after the proof of Lemma 6). It would be desirable to improve this bound. Our conjectural large sieve inequality would lead to a level of distribution $1/2-\varepsilon$.
@@ 2.1. Hecke characters for real quadratic fields
In this subsection, we provide an explicit description of Hecke characters for real quadratic fields, based on our more detailed description in [2, section 2], which in turn was based on material in [6], [7], [14, section 3.3], and [17]. Hecke characters are characters of the group of fractional ideals of $\mathbb{K}$ which are coprime to some modulus. We will use them to detect elements of $\mathcal{O}_{\mathbb{K}}$ in intersections of sectors and arithmetic progressions. To describe them properly, we set up some notation below.
@@
when $\alpha\in\mathbb{K}$ is such that $(\alpha)$ is coprime to $\mathfrak{q}$. If $(\alpha)$ is not coprime to $\mathfrak{q}$, then we set $\chi(\alpha):=0$. It is easy to see that $\chi(\alpha)$ is well-defined. We refer to $\chi$ as a Dirichlet character modulo $\mathfrak{q}$ as well. On the group $\mathcal{P}^{\text{frac}}_{\mathbb{K}}$ of principal fractional ideals of $\mathbb{K}$, Hecke characters modulo $\mathfrak{q}$ take the form
@@
where $\chi$ is a Dirichlet character modulo $\mathfrak{q}$ and $\chi_{\infty}$ is referred to as infinite part of $\chi_{\text{Hecke}}$. Every Hecke character $\chi_{\text{Hecke}}$ on $\mathcal{P}^{\text{frac}}_{\mathbb{K}}$ extends to $h$ distinct Hecke characters on the full group of fractional ideals $\mathcal{I}^{\text{frac}}_{\mathbb{K}}$. By convention, given a Hecke character $\chi_{\text{Hecke}}$ on $\mathcal{P}^{\text{frac}}_{\mathbb{K}}$, we will fix one of these extended characters and denote it by $\chi_{\text{Hecke}}$ as well. We say that the Hecke character $\chi_{\text{Hecke}}$ belongs to the Dirichlet character $\chi$. (More precisely, $\chi_{\text{Hecke}}$ should be viewed as a Hecke character to a modulus with finite part $\mathfrak{q}$ and infinite part $\mathfrak{m}_{\infty}$. For simplicity, we surpress $\mathfrak{m}_{\infty}$ throughout this article.) For $\chi_{\text{Hecke}}((\alpha))$ to be well-defined, it is necessary that
@@ 2.2. Sums with Hecke Characters
Let $X\geq 1$ and $\mathbb{K}$ be a number field of degree $n$ over $\mathbb{Q}$. Let $\chi_{\text{\rm Hecke}}$ be a non-trivial Hecke character modulo an ideal $\mathfrak{q}\in\mathcal{P}_{\mathbb{K}}$. Then
@@
Let $X\geq 2$, $A>0$ and $\mathbb{K}$ be a number field. Then there exists a positive constant $C=C(A,\mathbb{K})$ such that for any primitive Hecke character $\chi_{\text{\rm Hecke}}$ modulo an ideal $\mathfrak{q}\in\mathcal{P}_{\mathbb{K}}$ with $\mathcal{N}(\mathfrak{q})\leq\log^{A}X$, we have